OpenStudy (anonymous):

Use the substitution method to solve the following system of linear equations: 2x + 2y = 10 x = –y + 5 A. The solution is (4, –1). B. There is no solution. C. There are an infinite number of solutions. D. The solution is (2, –3).

OpenStudy (anonymous):

@iPwnBunnies @Luigi0210 @wolfe8 @phi @Whitemonsterbunny17

OpenStudy (ipwnbunnies):

Add 'y' to both sides of the second equation, and you'll see. Plus, you can factor something. c:

OpenStudy (anonymous):

Is it D?

OpenStudy (anonymous):

wait.... A?

OpenStudy (anonymous):

@dan815

OpenStudy (ipwnbunnies):

What did you find out when you added 'y' to both sides? .-.

OpenStudy (anonymous):

Idk! >.< I dont understand how to do this at all

OpenStudy (ipwnbunnies):

.-. Just do what I say, add 'y' to both sides in the second equation. 2x + 2y = 10 x + y = 5 Can you factor something out of the top equation?

OpenStudy (anonymous):

um... 10?

OpenStudy (anonymous):

or is it 5?

OpenStudy (ipwnbunnies):

._. What factor do 2x and 2y share?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

you multiply them by 2?

OpenStudy (ipwnbunnies):

Right! :D 2x + 2y = 10 2(x+y) = 10 Can't we also factor out a 2 out of the 10? 2(x+y) = 2(5)

OpenStudy (anonymous):

sorry im back and im kind of confused what to do next

OpenStudy (ipwnbunnies):

Well, you have a 2 factor on both sides. You can cancel them out. ._. You're left with x + y = 5 The same as the second equation. The first equation is just a multiple of the second equation. They would graph the same line. c:

OpenStudy (anonymous):

oh, ok so C

OpenStudy (ipwnbunnies):

Yep.

OpenStudy (anonymous):

Thanks dude! Really helped me out =D got an 88% :D